OpenStudy (anonymous):

Kate made a candle for her sister's birthday. The candle is shaped like an equilateral triangular pyramid. Each triangle has a base of 8 cm and a height that is approximately 6.9 cm. Kate wants to wrap the candle in plastic wrap before putting it into the gift bag. Find how much plastic wrap Kate will need to cover the candle. A. 220.8 cm2 B. 82.8 cm2 C.55.2 cm2 D.110.4 cm2

OpenStudy (anonymous):

a equilateral triangular pyramid has 4 triangular faces we need to find the area surface of the four faces area of one face=(1/2)bh=(1/2)8*6.9=27.6 sq cm surface area= 27.6*4=110.4 cm2